### v2

# Smile asymptotic for Bachelier Implied Volatility

Abstract
We investigate the asymptotic behaviour of the Implied Volatility in the Bachelier setting, extending the framework introduced by [4] for the Black–Scholes setting. Exploiting the theory of regular variation, we derive explicit expressions for the Bachelier Implied Volatility in the wings of the smile, linking these to the tail behaviour of the underlying’s returns’ distribution. Furthermore, we establish a direct connection between the analyticity strip of the returns’ characteristic function and the asymptotic formula for the Implied Volatility smile at extreme moneyness.
$(\ddagger)$ Politecnico di Milano, Department of Mathematics, 32 p.zza L. da Vinci, Milano
Keywords: volatility surface, Bachelier model, Lee’s moment formula.
MSC2020 subject classifications: 60E10, 91G20.
Address for correspondence:
Roberto Baviera
Department of Mathematics
Politecnico di Milano
32 p.zza Leonardo da Vinci
I-20133 Milano, Italy
Tel. +39-02-2399 4575
roberto.baviera@polimi.it
Smile asymptotic for Bachelier Implied Volatility
1 Introduction
In recent years, the Bachelier model [2, 10] has regained relevance in financial modelling, particularly in markets where the underlying can take negative values, such as interest rates and certain commodity futures [8]. Given this renewed interest, a refined understanding of the Implied Volatility within the Bachelier setting has become increasingly relevant.
In the context of Implied Volatility asymptotic, the work of Benaim and Friz [4] has significantly advanced the theoretical understanding of smile behaviour in the Black–Scholes framework. By leveraging the theory of regular variation [5], they established a rigorous connection between the tail behaviour of the underlying’s distribution and the wings of the Implied Volatility surface. Their framework provided a powerful extension of Lee’s moment formula [14], yielding precise asymptotic under mild integrability conditions, that are in fact satisfied in most models of practical interest. A detailed description of this problem can be found in excellent textbooks (see, e.g., [11, Ch.7]).
The first contribution of this letter is the extension of the result in [4] to the Bachelier setting. In particular, we investigate the asymptotic behaviour of the Bachelier Implied Volatility for large moneyness. We show that, under appropriate assumptions on the distribution of underlying’s returns, the wings of the Bachelier Implied Volatility can be characterized through explicit formulas.
Another contribution of this work is to establish a direct connection between the asymptotic behaviour of the Implied Volatility smile and the analyticity properties of the characteristic function of returns. In particular, following [3], that related the behaviour of the tails of a probability distribution respectively to the corresponding momenting generating function, we show that, under the appropriate condition, the boundaries of the analyticity strip in the complex plane determine the exponential decay rate of returns’ distribution’s tails, which are related to the behaviour of Bachelier Implied Volatility for large absolute values of moneyness.
The rest of the paper is organized as follows. In Section 2, we recall the Bachelier formula for European options and the theory of regular varying functions. In Section 3, we obtain the asymptotic behaviour of the Bachelier Implied Volatility for large strikes. Finally, we state our conclusions in Section 4.
2 Bachelier formula and regular varying functions
Let us consider a call option with moneyness $\kappa$  [^1] In this framework, if $F_{0}$ and $K$ are, respectively, the value of the underling at time 0 and option’s strike, and $t>0$ is option’s time-to-maturity, the moneyness $\kappa$ is defined as $\frac{K-F_{0}}{\sqrt{t}}$.. If we indicate with $\zeta$ the risk neutral returns [^2] In this framework, if $F_{0}$ and $F_{t}$ are, respectively, the value of the underling at time 0 and $t$, the return $\zeta$ is defined as $\frac{F_{t}-F_{0}}{\sqrt{t}}$., $F(\bullet)$ the cdf of $\zeta$, $f(\bullet)$ its pdf, and $\bar{F}(\bullet):=1-F(\bullet)$, then its undiscounted call price can be written as
$c(\kappa):=\mathbb{E}\left[\zeta-\kappa\right]^{+}=\int_{\kappa}^{+\infty}(\zeta-\kappa)\mathrm{d}F(\zeta)\,\,.$ (1)
The normalized Bachelier call price $c_{b}$ corresponds to the case with $\zeta$ distributed as a zero mean Gaussian rv with variance $\sigma^{2}$ (see, e.g., [7, eq.(3)])
$c_{b}(\kappa,\sigma):=-\kappa\,\Phi\left(-\frac{\kappa}{\sigma}\right)+\sigma\,\varphi\left(-\frac{\kappa}{\sigma}\right)\,\,$, (2)
with $\varphi(\bullet)$ and $\Phi(\bullet)$, respectively, the pdf and the cdf of a standard normal rv.
The Implied Volatility is the (unique) value [^3] The existence of a unique Implied Volatility for the Bachelier formula is a consequence of the continuity and monotonicity in $\sigma>0$ of the function $\sigma\mapsto c_{b}(\kappa,\sigma):\mathbb{R}^{+}\setminus\{0\}\to\mathbb{R}^{+}\setminus\{0\}$. $\mathrm{I}(\kappa)$ such that
$c_{b}\left(\kappa,\mathrm{I}(\kappa)\right)=c(\kappa)\,\,.$ (3)
This setting is an extension to real valued underling of [4], who consider the relation with Black-Scholes formula.
Similarly to (2), the normalized Bachelier put price is
$p_{b}(\kappa,\sigma):=\kappa\left(1-\,\Phi\left(-\frac{\kappa}{\sigma}\right)\right)+\sigma\varphi\left(-\frac{\kappa}{\sigma}\right)\,\,$, (4)
that can be written as in (3) in the following form
$p_{b}\left(\kappa,\mathrm{I}(\kappa)\right)=p(\kappa):=\int_{-\infty}^{\kappa}(\kappa-\zeta)\mathrm{d}F(\zeta)\,\,.$ (5)
The following Lemma identifies upper and lower bounds for the normalized Bachelier call formula.
Lemma 2.1.
(Bachelier formula bounds)
For $\beta>0$, $\forall\,y>0$
$e^{-\frac{1}{2\beta}y}g_{l}(y)\leq c_{b}\left(y,\sqrt{\beta y}\right)\leq e^{-\frac{1}{2\beta}y}g_{u}(y)\,\,$,
where $g_{l}(y)$ and $g_{u}(y)$ are positive and go to $\infty$ as $y\to+\infty$.
Proof.
Utilizing the (normalized) Bachelier formula (2), we have
$c_{b}\left(y,\sqrt{\beta y}\right)=\displaystyle-y\,\Phi\left(-\sqrt{\frac{y}{\beta}}\right)+\sqrt{\beta\,y}\,\varphi\left(-\sqrt{\frac{y}{\beta}}\right)\leq\sqrt{\frac{\beta\,y}{2\pi}}e^{-\frac{1}{2\beta}y}\,\,$,
where the inequality holds because the first addend is negative.
This proves the upper bound choosing
$g_{u}(y):=\sqrt{\frac{\beta\,y}{2\pi}}\,\,.$
For the lower bound, we recall that, for $x>0$, the following inequality holds (see, e.g., [16, Sec.2.26])
$\Phi(-x)\leq\frac{1}{x+\sqrt{x^{2}+\frac{4}{\pi}}}\varphi(-x)\,\,$,
thus,
$c_{b}\left(y,\sqrt{\beta y}\right)\geq e^{-\frac{1}{2\beta}y}\frac{1}{\sqrt{2\pi}}\left\{-\frac{y}{\sqrt{\frac{y}{\beta}}+\sqrt{\frac{y}{\beta}+\frac{4}{\pi}}}+\sqrt{\beta\,y}\right\}=g_{l}(y)e^{-\frac{1}{2\beta}y}\,\,$, (6)
with
$g_{l}(y):=\frac{\sqrt{\beta y}\sqrt{\frac{y}{\beta}+\frac{4}{\pi}}}{\sqrt{\frac{y}{\beta}}+\sqrt{\frac{y}{\beta}+\frac{4}{\pi}}}$
We recall the notion of regularly varying functions. We provide the formal definition and state a result from [5] concerning their asymptotic behaviour.
We give the following definition of regular varying functions (cf., [5, p.18]).
Definition 2.2.
A positive real-valued measurable function $g$ is regular varying with index $\theta\geq 0$, in symbols $g\in\mathscr{R}_{\theta}$, if $\forall\,\mu>0$
$\lim_{x\to+\infty}\frac{g(\mu x)}{g(x)}=\mu^{\theta}\,\,.$
[…]
The following result can be found in [5].
Theorem 2.3 (cf., [5, Th.4.12.10, p.255]).
Let $g\in\mathscr{R}_{\theta}$. Then, for $x\to+\infty$
$\ln\int_{x}^{+\infty}e^{-g(y)}\mathrm{d}y\sim g(x)$
3 Smile asymptotic
Let us introduce the following integrability conditions for the right (IR) and left (IL) tail of returns’ distribution
$\begin{cases}\text{(IR)}&\quad\exists\,\epsilon>0\quad\text{such that}\quad\mathbb{E}\left[e^{\epsilon\zeta}\right]<+\infty\\ \text{(IL)}&\quad\exists\,\epsilon>0\quad\text{such that}\quad\mathbb{E}\left[e^{-\epsilon\zeta}\right]<+\infty\end{cases}\,\,.$
The conditions above can also be expressed in terms of the analyticity properties of the characteristic function of the rv $\zeta$
$\phi(\xi):=\mathbb{E}\left[e^{i\xi\zeta}\right]=\int_{-\infty}^{+\infty}e^{i\xi\zeta}\mathrm{d}F(\zeta)\,\,.$
[…]
Moreover, if $\phi(\xi)$ is analytic in a neighbourhood of the origin, then, thanks to [15, Th.3.1, p.12], $\phi(\xi)$ is analytic in the whole horizontal strip with $\Im(\xi)\in(-\lambda_{-},\lambda_{+})$.
We demonstrate a first result, that relates the call price to the returns’ distribution.
Lemma 3.1.
Assume (IR), then
$c(\kappa)=\int_{\kappa}^{+\infty}\bar{F}(\zeta)\mathrm{d}\zeta\,\,.$
Proof.
Note that
$\bar{F}(\kappa)=\mathbb{P}\left(\zeta>\kappa\right)\leq e^{-\epsilon\kappa}\mathbb{E}\left[e^{\epsilon\zeta}\right]\,\,$, (7)
thus, $\bar{F}$ goes to 0 for $\kappa\to+\infty$ faster than exponentially.
The above property and integration by parts give
$c(\kappa):=$ $\int_{\kappa}^{+\infty}(\zeta-\kappa)\mathrm{d}F(\zeta)=-\int_{\kappa}^{+\infty}(\zeta-\kappa)\mathrm{d}\bar{F}(\zeta)=$
$=$ $-\left[(\zeta-\kappa)\bar{F}(\zeta)\right]_{\kappa}^{+\infty}+\int_{\kappa}^{+\infty}\bar{F}(\zeta)\mathrm{d}\zeta=\int_{\kappa}^{+\infty}\bar{F}(\zeta)\mathrm{d}\zeta$
We can introduce the following result about the Implied Volatility smile asymptotic, that is an extension to the Bachelier framework of Theorem 1.1 in [4]. Before proving the Theorem, we demonstrate a technical Lemma.
Lemma 3.2.
Assume (IR), then the following asymptotic properties of the Implied Volatility $\mathrm{I}(\kappa)$ hold
$\text{i)}\quad\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}=\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}=\frac{1}{2}\;\;,$
$\text{ii)}\quad\displaystyle\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\quad\text{as}\;\;\kappa\to+\infty\;\;.$
Proof.
Thanks to (7), condition (IR) implies that
$c(\kappa)\leq O\left(e^{-\epsilon\kappa}\right)\,\,$,
[…]
$\frac{c_{b}(\kappa,\mathrm{I}(\kappa))}{c_{b}(\kappa,\sqrt{\beta\,\kappa})}\leq\frac{O(e^{-\epsilon\,\kappa})}{g_{l}(\kappa)e^{-\frac{1}{2\beta}\kappa}}\rightarrow 0^{+}\quad\text{as}\quad\kappa\rightarrow+\infty\,\,.$ (8)
This means that, $\forall\,\epsilon>0$ such that (IR) holds, choosing $\beta=\frac{1}{2\epsilon}$ in (8), $\mathrm{I}(\kappa)$ is slower than $\sqrt{\frac{\kappa}{2\epsilon}}$, i.e. $\sqrt{\frac{\kappa}{2\epsilon}}=O(\mathrm{I}(\kappa))$, thus
$\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}=\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}=\frac{1}{2}\;\;.$
Moreover,
$\lim_{\kappa\to+\infty}\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}\,\left|\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)\right|}=+\infty\,\,$, (9)
that holds since, if $\mathrm{I}(\kappa)\not\to+\infty$, then the denominator in (9) does not diverge, otherwise, if $\mathrm{I}(\kappa)\to+\infty$, then $\mathrm{I}(\kappa)^{2}\,\ln\left(\mathrm{I}(\kappa)\right)=o\left(\mathrm{I}(\kappa)^{3}\right)=o\left(\kappa^{2}\right)$.
Equation (9) implies that, as $\kappa\to+\infty$
$\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)$
In the following Theorem, we don’t show a first implication, that, for $\theta\geq 1$,
$\ln f(\kappa)\in\mathscr{R}_{\theta}\quad\Rightarrow\quad-\ln\bar{F}(\kappa)\in\mathscr{R}_{\theta}\,\,$,
that has been proven in [4, Th.1.1, p.2].
Theorem 3.3 (Right wing formula).
Assume $\theta\geq 1$ and (IR). Then, for $\kappa\to+\infty$, the following implications hold
$\vskip 5.69046pt\text{i)}\;\;-\ln\bar{F}(\kappa)\in\mathscr{R}_{\theta}\;\;\Rightarrow\;\;\text{ii)}\;\;\begin{cases}-\ln c(\kappa)\sim-\ln\bar{F}(\kappa)&\\ -\ln c(\kappa)\in\mathscr{R}_{\theta}&\end{cases}\hskip-14.22636pt\;\;\Rightarrow\;\;\text{iii)}\;\;\frac{\mathrm{I}(\kappa)^{2}}{\kappa}\sim-\frac{\kappa}{2\ln\bar{F}(\kappa)}\,\,.$
Proof.
We prove that i) $\Rightarrow$ ii). Thanks to Lemma 3.1, we can prove that
$c(\kappa)=\int_{\kappa}^{+\infty}\bar{F}(\zeta)\mathrm{d}\zeta=\int_{\kappa}^{+\infty}e^{-\left(-\ln\bar{F}(\zeta)\right)}\mathrm{d}\zeta\,\,$,
thus, thanks to Theorem 2.3, with $g(\bullet)=-\ln\bar{F}(\bullet)$
$\ln c(\kappa)=-\ln\int_{\kappa}^{+\infty}e^{-\left(-\ln\bar{F}(\zeta)\right)}\mathrm{d}\zeta\sim-\ln\bar{F}(\kappa)\in\mathscr{R}_{\theta}\,\,.$ (10)
Let us prove that ii) $\Rightarrow$ iii). We first recall the following bounds for the function $\Phi(-x)$, for $x>0$ (see, e.g., [16, Sec.2.26])
$\varphi(-x)\frac{1}{x+\sqrt{x^{2}+2}}\leq\Phi(-x)\leq\varphi(-x)\frac{1}{x+\sqrt{x^{2}+\frac{4}{\pi}}}\,\,$,
from these inequalities, we obtain the following bounds for the normalized Bachelier formula (2) for any positive $\kappa$
$\varphi\left(-d\right)\left\{\sigma-\frac{\kappa}{d+\sqrt{d^{2}+\frac{4}{\pi}}}\right\}\leq c_{b}(\kappa,\sigma)\leq\varphi\left(-d\right)\left\{\sigma-\frac{\kappa}{d+\sqrt{d^{2}+2}}\right\}\,\,$,
with $d=\frac{\kappa}{\sigma}$. Thus, considering the Implied Volatility $I(\kappa)$ in (3) and $d(\kappa)=\frac{\kappa}{\mathrm{I}(\kappa)}$, we obtain
$\varphi\left(-d(\kappa)\right)\left\{\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+\frac{4}{\pi}}}\right\}\leq c(\kappa)\leq\varphi\left(-d(\kappa)\right)\left\{\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+2}}\right\}\,\,.$ (11)
Let us observe that
$\begin{cases}\displaystyle\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+\frac{4}{\pi}}}&=\displaystyle\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\vskip 8.53581pt\\ \vskip 8.53581pt\displaystyle\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+2}}&=\displaystyle\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\par\end{cases}\;\;.$ (12)
If we substitute (12) in (11) and apply the natural logarithm, we get
$-\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}-\ln\left(\sqrt{2\pi}\right)$ $+\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\right)\leq\ln c(\kappa)\leq$
$\leq\displaystyle-\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}-\ln\left(\sqrt{2\pi}\right)+\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\right)$
$\Leftrightarrow\,\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\right)\leq\ln c(\kappa)+\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}+\ln{\sqrt{2\pi}}\leq\displaystyle\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\right)\,\,.$ (13)
We define
$\begin{cases}\displaystyle\varepsilon_{1}(\kappa):=&\ln c(\kappa)+\displaystyle\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\\[8.53581pt] \displaystyle\varepsilon_{2}(\kappa):=&\varepsilon_{1}(\kappa)+\ln{\sqrt{2\pi}}\end{cases}\,\,.$
Observing that $\varepsilon_{1}(\kappa)=o\left(\displaystyle\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)$ concludes the proof. Indeed, if $\varepsilon_{1}(\kappa)=o\left(\displaystyle\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)$ as $\kappa\to+\infty$, then
$\ln c(\kappa)=-\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}+o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\quad\Leftrightarrow\quad\frac{\ln c(\kappa)}{\kappa}=-\frac{\kappa}{2\,\mathrm{I}(\kappa)^{2}}+o\left(\frac{\kappa}{2\,\mathrm{I}(\kappa)^{2}}\right)\,\,$,
this proves that, as $\kappa\to+\infty$,
$\frac{\ln c(\kappa)}{\kappa}\sim-\frac{\kappa}{2\,\mathrm{I}(\kappa)^{2}}\quad\Leftrightarrow\quad\frac{\mathrm{I}(\kappa)^{2}}{\kappa}\sim-\frac{\kappa}{2\,\ln c(\kappa)}\sim-\frac{\kappa}{2\,\ln\bar{F}(\kappa)}\,\,$,
where the last asymptotic holds thanks to (10).
To conclude the proof, we demonstrate that $\varepsilon_{1}(x)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)$ as $\kappa\to+\infty$. From (13), we have
$\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\right)\leq\varepsilon_{2}(\kappa)\leq\displaystyle\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\right)\,\,$, (14)
applying point i) of Lemma 3.2, the inequalities above imply that
$\varepsilon_{2}(\kappa)\sim\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)\,\,.$ (15)
Finally, the asymptotic in (15) and point ii) of Lemma 3.2 imply that
$\varepsilon_{2}(\kappa)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\Rightarrow\varepsilon_{1}(\kappa)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\quad\text{as}\;\,\kappa\to+\infty$
In the Theorem above, we have showed the relation between the Implied Volatility for large positive moneyness and returns’ distribution. Now, in the same way, we can establish a connection between the Implied Volatility for large negative moneyness and returns’ distribution.
We first show the relation between the put price and the returns’ cdf. After that, we prove the formula for the Implied Volatility for large negative moneyness, omitting, as in Theorem 3.3, the implication that, for $\theta\geq 1$,
$\ln f(-\kappa)\in\mathscr{R}_{\theta}\quad\Rightarrow\quad-\ln F(-\kappa)\in\mathscr{R}_{\theta}\,\,.$
Lemma 3.4.
Assume (IL), then
$p(\kappa)=\int_{-\infty}^{\kappa}F(\zeta)\mathrm{d}\zeta\,\,.$
Proof.
The proof is the analogous of Lemma 3.1
Lemma 3.5.
Assume (IL), then the following asymptotic properties of the Implied Volatility $\mathrm{I}(\kappa)$ hold
$\text{i)}\quad\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(-\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(-\kappa)^{2}}}=\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(-\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(-\kappa)^{2}}}=\frac{1}{2}\;\;,$
$\text{ii)}\quad\displaystyle\ln\left(\mathrm{I}(-\kappa)\right)+\ln\left(\frac{1}{2}\right)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(-\kappa)^{2}}\right)\quad\text{as}\;\;\kappa\to+\infty\;\;.$
Proof.
The proof is the analogous of Lemma 3.2, considering that, $\forall\,\sigma>0$ and $\kappa\in\mathbb{R}$
$p_{b}(-\kappa,\sigma)=c_{b}(\kappa,\sigma)\;\;,$
and that condition (IL) inplies that
$p(-\kappa)\leq O(e^{-\epsilon\kappa})\quad\text{as}\;\;\kappa\to+\infty$
Theorem 3.6 (Left wing formula).
Assume $\theta\geq 1$ and (IL). Then, for $\kappa\to+\infty$, the following implications hold
$\text{i)}\;\;-\ln F(-\kappa)\in\mathscr{R}_{\theta}\,\,\Rightarrow\,\,\text{ii)}\;\;\begin{cases}-\ln p(-\kappa)\sim-\ln F(-\kappa)&\\ -\ln p(-\kappa)\in\mathscr{R}_{\theta}&\end{cases}\hskip-14.22636pt\,\,\Rightarrow\,\,\text{iii)}\;\;\frac{\mathrm{I}(-\kappa)^{2}}{\kappa}\sim-\frac{\kappa}{2\ln F(-\kappa)}\,\,.$
Proof.
The proof is the analogous of Theorem 3.3, using Lemma 3.4 and 3.5
As a consequence of Theorem 3.3 and 3.6, it holds that if $\ln\bar{F}(\kappa)$ and $\ln F(-\kappa)\in\mathscr{R}_{\theta}$, then
$\frac{\mathrm{I}(\kappa)^{2}}{\kappa}\sim\begin{cases}&\displaystyle-\frac{\kappa}{2\,\ln\bar{F}(\kappa)}\quad\text{as}\;\,\kappa\to+\infty\,\,,\\[14.22636pt] &\displaystyle-\frac{\kappa}{2\,\ln F(\kappa)}\quad\text{as}\;\,\kappa\to-\infty\end{cases}\,\,.$
For most models, the asymptotic behaviour of the returns’ distribution is not known explicitly, but we can relate it, and then our formula, to the behaviour of the characteristic function of the rv $\zeta$ at the boundaries of the analyticity strip.
We now introduce the following condition.
Condition I: Let $\phi(\xi)$ analytic in the strip $\Im(\xi)\in(-\lambda_{-},\lambda_{+})$ with $\lambda_{\pm}>0$ and finite. Suppose that for some $n\geq 0$,
$(-i)^{n}\phi^{(n)}\left(\pm i\left(\lambda_{\pm}-s\right)\right)\sim s^{-\rho_{\pm}}\,l_{\pm}\left(\frac{1}{s}\right)\quad\text{as $s\to 0^{+}$}\;\;,$
[…]
This assumption ensures that some derivative of the characteristic function diverges at $\pm\lambda_{\pm}$ in a regularly varying manner.
We state our result in the following Theorem.
Theorem 3.7.
Consider $\lambda_{\pm}>0$ finite. Then, the following implications hold:
$\text{i)}\quad\text{Condition I}\;\Rightarrow\;\text{ii)}\quad\begin{cases}\ln\bar{F}(\kappa)\sim-\lambda_{-}\kappa,\\ \ln F(-\kappa)\sim-\lambda_{+}\kappa\end{cases}\;\;\text{as }\kappa\to+\infty\;\Rightarrow\;\text{iii)}\quad\lim_{\kappa\to\pm\infty}\frac{\mathrm{I}(\kappa)^{2}}{|\kappa|}=\frac{1}{2\lambda_{\mp}}.$
Proof.
We prove that i) $\Rightarrow$ ii). Let us consider the moment generating function of the random variable $\zeta$:
$M(s):=\mathbb{E}\left[e^{s\zeta}\right]=\phi(-is),\quad s\in(-\lambda_{+},\lambda_{-})\,\,.$
Condition I is equivalent to
$M^{(n)}\left(\mp\left(\lambda_{\pm}-s\right)\right)\sim s^{-\rho_{\pm}}\,l_{\pm}\left(\frac{1}{s}\right)\quad\text{as $s\to 0^{+}$}\;\;,$
for some $\rho_{\pm}>0$ and $l_{\pm}\in\mathscr{R}_{0}$, where $M^{(n)}$ denotes the $n$-th derivative of the moment generating function. Indeed
$M^{(n)}\left(\mp\left(\lambda_{\pm}-s\right)\right)=(-i)^{n}\phi^{(n)}\left(\pm i\left(\lambda_{\pm}-s\right)\right)\;\;.$
Thus, we can apply Criterion $1$ in [3, Criterion 1, p.20], and conclude that
$\begin{cases}\ln\bar{F}(\kappa)\sim-\lambda_{-}\kappa,\\ \ln F(-\kappa)\sim-\lambda_{+}\kappa\end{cases}\quad\text{as}\;\;\kappa\to+\infty\;\;.$
Finally, we prove that ii) $\Rightarrow$ iii). Thanks to the fact that the function $\phi(\xi)$ is analytic in the strip $\Im(\xi)\in(-\lambda_{-},\lambda_{+})$, then (IR) and (IL) hold. Moreover, thanks to ii), $\ln\bar{F}(\kappa)$ and $\ln F(-\kappa)\in\mathscr{R}_{1}$, and we can apply Theorem 3.3 and 3.6
$\lim_{\kappa\to+\infty}\frac{\mathrm{I}(\kappa)^{2}}{\kappa}$ $=-\lim_{\kappa\to+\infty}\frac{\kappa}{2\,\ln\bar{F}(\kappa)}=\frac{1}{2\,\lambda_{-}}\,\,$,
$\lim_{\kappa\to-\infty}\frac{\mathrm{I}(\kappa)^{2}}{\kappa}$ $=-\lim_{\kappa\to-\infty}\frac{\kappa}{2\,\ln{F}(\kappa)}=-\frac{1}{2\,\lambda_{+}}$
Condition I is fairly general and is satisfied by most models commonly used to describe returns in quantitative finance, where the characteristic function is analytic in the whole horizontal strip $\Im(\xi)\in(-\lambda_{-},\lambda_{+})$. This includes stochastic volatility models (e.g., Heston model [13]), Lévy models (such as Meixner [12], Tempered Stable, NTS (see, e.g., [9, Ch.4])), and Additive models (e.g., [6, 1]). [^9] Some models satisfy an even stronger condition, where the singularities for $\xi=\pm\,i\,\lambda_{\pm}$ are poles, as is the case for the Heston model and the Meixner process.
For models that generalise the Bachelier framework –where returns are linearly related to the underlying asset price, as in (1)– Theorem 3.7 implies that the squared Implied Volatility grows linearly for large absolute values of moneyness $\kappa$, with a slope inversely proportional to the boundaries of the analyticity strip of the characteristic function. [^10] Theorem 3.7 is even more general. Its first implication can be applied to models –commonly used in quantitative finance– defined on log-returns, which yield Black-Scholes-type dynamics, such as the Heston model or exponential Lévy models [cf., 9, Ch.8].
4 Conclusions
In this letter, we have extended the asymptotic analysis of the Implied Volatility smile from the Black–Scholes setting, as developed in [4], to the Bachelier model, that is particularly relevant in markets where the underlying asset can take negative values.
Using the theory of regular variation, we derived explicit asymptotic expressions for the Bachelier Implied Volatility in the wings of the volatility smile, i.e., for large positive or negative moneyness. We have proven that there exists a direct link between the tail behaviour of the distribution of the underling’s returns and the corresponding Implied Volatility behaviour.
Moreover, in cases where the distribution of returns is not available in closed form but the characteristic function is known, we showed that the asymptotic behaviour of the Implied Volatility can still be characterised by our formula. In particular, we established that the boundaries of the analyticity strip of returns’ characteristic function directly determines the behaviour of Bachelier Implied Volatility in the wings of the smile for all models used in quantitative finance.