### v3

# Smile asymptotics for Bachelier implied volatility

Abstract
We investigate the asymptotic behaviour of the implied volatility in the Bachelier setting, extending the large-strike results established for the Black-Scholes framework. Exploiting the theory of regular variation, we derive explicit expressions for the Bachelier implied volatility in the wings of the smile, directly linking them to the tail decay of the underlying returns’ distribution. Furthermore, we establish a rigorous connection between the analyticity strip of the characteristic function and the asymptotic slope of the volatility smile. Moreover, we show that if the implied variance grows linearly for large absolute moneyness degree, the underlying returns’ distribution must exhibit exponential tail decay, and the corresponding characteristic function is analytic in a horizontal strip of the complex plane. These findings characterizes valid models based solely on the observable asymptotic behaviour of the smile.
$(\ddagger)$ Politecnico di Milano, Department of Mathematics, 32 p.zza L. da Vinci, Milano
Keywords: volatility surface, Bachelier model, Lee’s moment formula.
Address for correspondence:
Roberto Baviera
Department of Mathematics
Politecnico di Milano
32 p.zza Leonardo da Vinci
I-20133 Milano, Italy
Tel. +39-02-2399 4575
roberto.baviera@polimi.it
Smile asymptotics for Bachelier implied volatility
1 Introduction
In recent years, the Bachelier model (Bachelier 1900, Davis and Etheridge 2006) has regained relevance in financial modelling, particularly in markets where the underlying can take negative values, such as interest rates and certain commodity futures (CME 2020, ICE 2020). Given this renewed interest, a refined understanding of the implied volatility within the Bachelier setting has become increasingly relevant.
In the context of implied volatility asymptotics, the work of Benaim and Friz (2009) has significantly advanced the theoretical understanding of smile behaviour in the Black–Scholes framework. By leveraging the theory of regular variation (Bingham et al. 1987), they established a rigorous connection between the tail behaviour of the underlying’s distribution and the wings of the implied volatility surface. Their framework provided a powerful extension of Lee (2004) moment formula, yielding precise asymptotics under mild integrability conditions that are, in fact, satisfied by most models of practical interest. A detailed description of this problem can be found in excellent textbooks (see, e.g., Gatheral 2011, Ch.7).
Three are the main contributions of the paper.
The first contribution of this paper is to extend the result of Benaim and Friz (2009) to the Bachelier setting. In particular, we investigate the asymptotic behaviour of the Bachelier implied volatility for large moneyness degree. We show that, under appropriate assumptions on the distribution of underlying’s returns, the wings of the Bachelier implied volatility can be characterised through explicit formulas.
Another contribution of this work is to establish a direct connection between the asymptotic behaviour of the implied volatility smile and the analyticity properties of the characteristic function of returns. In particular, following Benaim and Friz (2008), that related the behaviour of the tails of a probability distribution to the corresponding moment generating function, we show that the boundaries of the analyticity strip in the complex plane determine the exponential decay rate of returns’ distribution tails, which are related to the behaviour of Bachelier implied volatility for large absolute values of moneyness degree.
The third contribution is to show that the asymptotic behaviour of the implied volatility determines the behaviour of the tails of the returns’ probability distribution. In particular, we prove that if the implied variance grows linearly in the absolute value of the moneyness degree, for large moneyness degree, then the returns’ probability distribution tails exhibit exponential decay, and the corresponding characteristic function is analytical in a horizontal strip in the complex plane.
The rest of the paper is organised as follows. In Section 2, we recall the Bachelier formula for European options and the theory of regular varying functions. In Section 3, we obtain the asymptotic behaviour of the Bachelier implied volatility for large strikes, and we show that if the implied variance is asymptotic to the moneyness degree, then the returns’ characteristic function is analytical in a horizontal strip of the complex plane. Finally, we state our conclusions in Section 4.
2 Bachelier formula and regular varying functions
Let us consider a call option with moneyness degree $\kappa$. [^1] In this framework, if $F_{0}$ and $K$ are, respectively, the value of the underling at time 0 and option’s strike, and $t>0$ is option’s time-to-maturity, the moneyness degree $\kappa$ is defined as $\frac{K-F_{0}}{\sqrt{t}}$. If we indicate with $\zeta$ the risk-neutral returns, [^2] In this framework, if $F_{0}$ and $F_{t}$ are, respectively, the value of the underlying at time 0 and $t$, the return $\zeta$ is defined as $\frac{F_{t}-F_{0}}{\sqrt{t}}$. $F(\bullet)$ the cdf of $\zeta$ and $\bar{F}(\bullet):=1-F(\bullet)$, then its undiscounted call price can be written as
$c(\kappa):=\mathbb{E}\left[\zeta-\kappa\right]^{+}=\int_{\kappa}^{+\infty}(\zeta-\kappa)\mathrm{d}F(\zeta)\,\,.$ (1)
The normalized Bachelier call price $c_{b}$ corresponds to the case with $\zeta$ distributed as a zero mean Gaussian rv with variance $\sigma^{2}$ (see, e.g., Choi et al. 2022, eq.(3))
$c_{b}(\kappa,\sigma):=-\kappa\,\Phi\left(-\frac{\kappa}{\sigma}\right)+\sigma\,\varphi\left(-\frac{\kappa}{\sigma}\right)\,\,$, (2)
with $\varphi(\bullet)$ and $\Phi(\bullet)$, respectively, the pdf and the cdf of a standard normal rv.
The implied volatility is the (unique) value  [^3] The existence of a unique implied volatility for the Bachelier formula is a consequence of the continuity and monotonicity in $\sigma>0$ of the function $\sigma\mapsto c_{b}(\kappa,\sigma):\mathbb{R}^{+}\setminus\{0\}\to\mathbb{R}^{+}\setminus\{0\}$. $\mathrm{I}(\kappa)$ such that
$c_{b}\left(\kappa,\mathrm{I}(\kappa)\right)=c(\kappa)\,\,.$ (3)
This setting is an extension to real valued underling of Benaim and Friz (2009), who consider the relation for the Black-Scholes formula.
Similarly to (2), the normalized Bachelier put price is
$p_{b}(\kappa,\sigma):=\kappa\left(1-\,\Phi\left(-\frac{\kappa}{\sigma}\right)\right)+\sigma\varphi\left(-\frac{\kappa}{\sigma}\right)\,\,$, (4)
that can be written as in (1) and (3) as follows
$p_{b}\left(\kappa,\mathrm{I}(\kappa)\right)=p(\kappa):=\int_{-\infty}^{\kappa}(\kappa-\zeta)\mathrm{d}F(\zeta)\,\,.$ (5)
The following Lemma identifies upper and lower bounds for the normalised Bachelier call formula.
Lemma 2.1.
(Bachelier formula bounds)
For any $\beta>0$, $y>0$
$e^{-\frac{1}{2\beta}y}g_{l}(y)\leq c_{b}\left(y,\sqrt{\beta y}\right)\leq e^{-\frac{1}{2\beta}y}g_{u}(y)\,\,$,
where $g_{l}(y)$ and $g_{u}(y)$ are positive and go to $\infty$ as $y\to+\infty$.
Proof.
Utilizing the (normalized) Bachelier formula (2), we have
$c_{b}\left(y,\sqrt{\beta y}\right)=\displaystyle-y\,\Phi\left(-\sqrt{\frac{y}{\beta}}\right)+\sqrt{\beta\,y}\,\varphi\left(-\sqrt{\frac{y}{\beta}}\right)\leq\sqrt{\frac{\beta\,y}{2\pi}}e^{-\frac{1}{2\beta}y}\,\,$,
where the inequality holds because the first addend is negative.
This proves the upper bound choosing
$g_{u}(y):=\sqrt{\frac{\beta\,y}{2\pi}}\,\,.$
For the lower bound, we recall that, for $x>0$, the following inequality holds (see, e.g., Mitrinovic 1970, Sec.2.26, p.179)
$\Phi(-x)\leq\frac{1}{x+\sqrt{x^{2}+\frac{4}{\pi}}}\varphi(-x)\,\,$,
thus,
$c_{b}\left(y,\sqrt{\beta y}\right)\geq e^{-\frac{1}{2\beta}y}\frac{1}{\sqrt{2\pi}}\left\{-\frac{y}{\sqrt{\frac{y}{\beta}}+\sqrt{\frac{y}{\beta}+\frac{4}{\pi}}}+\sqrt{\beta\,y}\right\}=g_{l}(y)e^{-\frac{1}{2\beta}y}\,\,$, (6)
with
$g_{l}(y):=\frac{\sqrt{\beta y}\sqrt{\frac{y}{\beta}+\frac{4}{\pi}}}{\sqrt{\frac{y}{\beta}}+\sqrt{\frac{y}{\beta}+\frac{4}{\pi}}}$
We recall the notion of regularly varying functions. We provide the formal definition (cf., Bingham et al. 1987, p.18) and state a result from Bingham et al. (1987) concerning their asymptotic behaviour.
Definition 2.2.
A positive real-valued measurable function $g$ is regular varying with index $\theta\geq 0$, in symbols $g\in\mathscr{R}_{\theta}$, if $\forall\,\mu>0$
$\lim_{x\to+\infty}\frac{g(\mu x)}{g(x)}=\mu^{\theta}\,\,.$
[…]
The following result can be found in Bingham et al. (1987, , Thm.4.12.10, p.255).
Theorem 2.3.
Let $g\in\mathscr{R}_{\theta}$. Then, for $x\to+\infty$
$\ln\int_{x}^{+\infty}e^{-g(y)}\mathrm{d}y\sim g(x)\,\,.$
3 Smile asymptotics
In this Section, we derive the main results of the paper concerning the asymptotic behaviour of the Bachelier implied volatility for large absolute values of moneyness degree. We establish a rigorous link between the decay rate of the underlying’s return distribution tails and the slope of the implied variance for large absolute moneyness degree.
We proceed in two steps. First, we investigate the direct problem: assuming a specific tail behaviour for the returns (expressed in terms of regular variation), we deduce the corresponding asymptotic formula for the implied volatility in the wings of the smile. Subsequently, we address the inverse problem, showing that the asymptotic behaviour of the implied volatility itself determines the analytic properties of the characteristic function of returns. This latter result is particularly significant as it allows for model characterisation based on observable market data.
We begin by presenting a preliminary result that relates the call and put prices to the cdf of returns.
Lemma 3.1.
The following holds for the call price $c(\kappa)$ in (1)
$c(\kappa)=\int_{\kappa}^{+\infty}\bar{F}(\zeta)\mathrm{d}\zeta\,\,$,
while for the put price $p(\kappa)$ in (5)
$p(\kappa)=\int_{-\infty}^{\kappa}F(\zeta)\mathrm{d}\zeta\,\,.$
Proof.
Let us recall that
$\zeta-\kappa]^{+}=\int_{\kappa}^{+\infty}\mathbbm{1}_{\{\zeta>x\}}\mathrm{d}x\,\,.$
Since the integrand is non-negative, by Fubini’s Theorem (cf. Rudin 1986, Thm.8.8, p.164), we can interchange the order of expectation and integration:
$c(\kappa)=\int_{\kappa}^{+\infty}\mathbb{E}\left[\mathbbm{1}_{\{\zeta>x\}}\right]\mathrm{d}x=\int_{\kappa}^{+\infty}\mathbb{P}(\zeta>x)\mathrm{d}x=\int_{\kappa}^{+\infty}\bar{F}(x)\mathrm{d}x\,\,$,
this proves the thesis. The proof for the put price is analogous ∎
We can introduce the following result on the asymptotic implied volatility smile, that is an extension to the Bachelier framework of Theorem 1.1 in Benaim and Friz (2009). Before proving the Theorem, we demonstrate two technical Lemmas.
Lemma 3.2.
(Right wing properties).
[…]
$\text{a)}\quad\displaystyle\exists\,\epsilon>0\quad\text{such that}\quad c(\kappa)\leq O\left(e^{-\epsilon\kappa}\right)\quad\text{as}\quad\kappa\to+\infty\;\;,$
$\text{b)}\quad\displaystyle\exists\,\mu>0\quad\text{such that}\quad\mathrm{I}(\kappa)\sim\mu\sqrt{\kappa}\quad\text{as}\quad\kappa\to+\infty\;\;,$
then the following asymptotic properties of the implied volatility $\mathrm{I}(\kappa)$ hold
$\text{i)}\quad\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}=\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}=\frac{1}{2}\;\;,$
$\text{ii)}\quad\displaystyle\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\quad\text{as}\;\;\kappa\to+\infty\;\;.$
Proof.
If a) holds, then, thanks to Lemma 2.1, $\forall\,\beta>0$ such that $\frac{1}{2\beta}\leq\epsilon$
$\frac{c_{b}(\kappa,\mathrm{I}(\kappa))}{c_{b}(\kappa,\sqrt{\beta\,\kappa})}\leq\frac{O(e^{-\epsilon\,\kappa})}{g_{l}(\kappa)e^{-\frac{1}{2\beta}\kappa}}\rightarrow 0^{+}\quad\text{as}\quad\kappa\rightarrow+\infty\,\,.$ (7)
This means that, $\forall\,\epsilon>0$ such that a) holds, choosing $\beta=\frac{1}{2\epsilon}$ in (7), there exists $\kappa_{0}$ such that for all $\kappa>\kappa_{0}$, $\mathrm{I}(\kappa)<\sqrt{\frac{\kappa}{2\epsilon}}$, thus
$\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(\kappa)^{2}}}=\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}=\frac{1}{2}\;\;,$
that proves i) if assumption a) holds. Moreover,
$\lim_{\kappa\to+\infty}\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}\,\left|\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)\right|}=+\infty\,\,$, (8)
that holds since, if $\mathrm{I}(\kappa)\not\to+\infty$, then the denominator in (8) does not diverge, otherwise, if $\mathrm{I}(\kappa)\to+\infty$, then $\mathrm{I}(\kappa)^{2}\,\ln\left(\mathrm{I}(\kappa)\right)=o\left(\mathrm{I}(\kappa)^{3}\right)=o\left(\kappa^{2}\right)$.
Equation (8) implies that, as $\kappa\to+\infty$
$\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\,\,$,
that proves ii) if assumption a) holds.
  Finally, under assumption b), properties i) and ii) follow straightforwardly from the asymptotic behaviour $\mathrm{I}(\kappa)\sim\mu\sqrt{\kappa}$ as $\kappa\to+\infty$, observing that $\mathrm{I}(\kappa)^{2}$ scales linearly with $\kappa$
Lemma 3.3.
(Left wing properties).
Assume that one of the following holds
$\text{a)}\quad\displaystyle\exists\,\epsilon>0\quad\text{such that}\quad p(-\kappa)\leq O\left(e^{-\epsilon\kappa}\right)\quad\text{as}\quad\kappa\to+\infty\;\;,$
$\text{b)}\quad\displaystyle\exists\,\mu>0\quad\text{such that}\quad I(-\kappa)\sim\mu\sqrt{\kappa}\quad\text{as}\quad\kappa\to+\infty\;\;,$
then the following asymptotic properties of the implied volatility $\mathrm{I}(\kappa)$ hold
$\text{i)}\quad\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(-\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\,\mathrm{I}(-\kappa)^{2}}}=\displaystyle\lim_{\kappa\to+\infty}\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(-\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(-\kappa)^{2}}}=\frac{1}{2}\;\;,$
$\text{ii)}\quad\displaystyle\ln\left(\mathrm{I}(-\kappa)\right)+\ln\left(\frac{1}{2}\right)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(-\kappa)^{2}}\right)\quad\text{as}\;\;\kappa\to+\infty\;\;.$
Proof.
The proof is analogous to Lemma 3.2, considering that, $\forall\,\sigma>0$ and $\kappa\in\mathbb{R}$
$p_{b}(-\kappa,\sigma)=c_{b}(\kappa,\sigma)$
The following Theorem extends to the Bachelier framework Theorem 1.1 in Benaim and Friz (2009), which has been proven in the Black-Scholes framework.
Theorem 3.4.
(Right wing formula)
Assume $\theta\geq 1$. Then, for $\kappa\to+\infty$, the following implications hold
$\vskip 5.69046pt\text{i)}\;\;-\ln\bar{F}(\kappa)\in\mathscr{R}_{\theta}\;\;\Rightarrow\;\;\text{ii)}\;\;\begin{cases}-\ln c(\kappa)\sim-\ln\bar{F}(\kappa)&\\ -\ln c(\kappa)\in\mathscr{R}_{\theta}&\end{cases}\hskip-14.22636pt\;\;\Rightarrow\;\;\text{iii)}\;\;\frac{\mathrm{I}(\kappa)^{2}}{\kappa}\sim-\frac{\kappa}{2\ln\bar{F}(\kappa)}\,\,.$
Proof.
We prove that i) $\Rightarrow$ ii). Thanks to Lemma 3.1,
$c(\kappa)=\int_{\kappa}^{+\infty}\bar{F}(\zeta)\mathrm{d}\zeta=\int_{\kappa}^{+\infty}e^{-\left(-\ln\bar{F}(\zeta)\right)}\mathrm{d}\zeta\,\,$,
thus, thanks to Theorem 2.3, with $g(\bullet)=-\ln\bar{F}(\bullet)$
$\ln c(\kappa)=-\ln\int_{\kappa}^{+\infty}e^{-\left(-\ln\bar{F}(\zeta)\right)}\mathrm{d}\zeta\sim-\ln\bar{F}(\kappa)\in\mathscr{R}_{\theta}\,\,.$ (9)
Let us prove that ii) $\Rightarrow$ iii). We first recall the following bounds for the function $\Phi(-x)$, for $x>0$ (see, e.g., Mitrinovic 1970, Sec.2.26, p.179)
$\varphi(-x)\frac{1}{x+\sqrt{x^{2}+2}}\leq\Phi(-x)\leq\varphi(-x)\frac{1}{x+\sqrt{x^{2}+\frac{4}{\pi}}}\,\,.$
From these inequalities, we obtain the following bounds for the normalized Bachelier formula (2) for any positive $\kappa$
$\varphi\left(-d\right)\left\{\sigma-\frac{\kappa}{d+\sqrt{d^{2}+\frac{4}{\pi}}}\right\}\leq c_{b}(\kappa,\sigma)\leq\varphi\left(-d\right)\left\{\sigma-\frac{\kappa}{d+\sqrt{d^{2}+2}}\right\}\,\,$,
with $d:=\frac{\kappa}{\sigma}$. Thus, considering the implied volatility $\mathrm{I}(\kappa)$ in (3) and $d(\kappa):=\frac{\kappa}{\mathrm{I}(\kappa)}$, we obtain
$\varphi\left(-d(\kappa)\right)\left\{\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+\frac{4}{\pi}}}\right\}\leq c(\kappa)\leq\varphi\left(-d(\kappa)\right)\left\{\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+2}}\right\}\,\,.$ (10)
Let us observe that
$\begin{cases}\displaystyle\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+\frac{4}{\pi}}}&=\displaystyle\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\vskip 8.53581pt\\ \vskip 8.53581pt\displaystyle\mathrm{I}(\kappa)-\frac{\kappa}{d(\kappa)+\sqrt{d(\kappa)^{2}+2}}&=\displaystyle\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\par\end{cases}\;\;.$ (11)
If we substitute (11) in (10) and apply the natural logarithm, we get
$-\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}-\ln\left(\sqrt{2\pi}\right)$ $+\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\right)\leq\ln c(\kappa)\leq$
$\leq\displaystyle-\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}-\ln\left(\sqrt{2\pi}\right)+\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\right)$
$\Leftrightarrow\,\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\right)\leq\ln c(\kappa)+\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}+\ln{\sqrt{2\pi}}\leq\displaystyle\ln\left(\frac{\mathrm{I}(\kappa)\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\right)\,\,.$ (12)
We define
$\begin{cases}\displaystyle\varepsilon_{1}(\kappa):=&\ln c(\kappa)+\displaystyle\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\\[8.53581pt] \displaystyle\varepsilon_{2}(\kappa):=&\varepsilon_{1}(\kappa)+\ln{\sqrt{2\pi}}\end{cases}\,\,.$ (13)
If $\varepsilon_{1}(\kappa)=o\left(\displaystyle\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)$ as $\kappa\to+\infty$, then
$\ln c(\kappa)=-\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}+o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\quad\Leftrightarrow\quad\frac{\ln c(\kappa)}{\kappa}=-\frac{\kappa}{2\,\mathrm{I}(\kappa)^{2}}+o\left(\frac{\kappa}{2\,\mathrm{I}(\kappa)^{2}}\right)\,\,$,
this proves that, as $\kappa\to+\infty$,
$\frac{\ln c(\kappa)}{\kappa}\sim-\frac{\kappa}{2\,\mathrm{I}(\kappa)^{2}}\quad\Leftrightarrow\quad\frac{\mathrm{I}(\kappa)^{2}}{\kappa}\sim-\frac{\kappa}{2\,\ln c(\kappa)}\sim-\frac{\kappa}{2\,\ln\bar{F}(\kappa)}\,\,$, (14)
where the last asymptotic holds thanks to (9).
To conclude the proof, we have to demonstrate that $\varepsilon_{1}(\kappa)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)$ as $\kappa\to+\infty$. From (12), we have
$\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+\frac{4}{\pi}\mathrm{I}(\kappa)^{2}}}\right)\leq\varepsilon_{2}(\kappa)\leq\displaystyle\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}{\kappa+\sqrt{\kappa^{2}+2\,\mathrm{I}(\kappa)^{2}}}\right)\,\,{\color[rgb]{0,0,1}.}$ (15)
We recall that if $\ln c(\kappa)\in\mathscr{R}_{\theta}$ with $\theta\geq 1$, then the call price $c(\kappa)$ decays at least exponentially as $\kappa\to+\infty$. Therefore, assumption a) of Lemma 3.2 is satisfied. Applying point i) of Lemma 3.2, the inequalities above imply that
$\varepsilon_{2}(\kappa)\sim\ln\left(\mathrm{I}(\kappa)\right)+\ln\left(\frac{1}{2}\right)\,\,.$ (16)
Finally, the asymptotic in (16) and point ii) of Lemma 3.2 imply that
$\varepsilon_{2}(\kappa)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\Rightarrow\varepsilon_{1}(\kappa)=o\left(\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\right)\quad\text{as}\;\,\kappa\to+\infty$
where the last implication is due to the definition of $\varepsilon_{2}(\kappa)$ in (13) ∎
In the Theorem above, we have shown the relation between the implied volatility for large positive moneyness degree and the returns’ distribution. Now, in the same way, we can establish a connection between the implied volatility for large negative moneyness degrees and returns’ distribution that extends Theorem 1.2 in Benaim and Friz (2009) to the Bachelier framework.
Theorem 3.5.
(Left wing formula)
Assume $\theta\geq 1$. Then, for $\kappa\to+\infty$, the following implications hold
$\text{i)}\;\;-\ln F(-\kappa)\in\mathscr{R}_{\theta}\,\,\Rightarrow\,\,\text{ii)}\;\;\begin{cases}-\ln p(-\kappa)\sim-\ln F(-\kappa)&\\ -\ln p(-\kappa)\in\mathscr{R}_{\theta}&\end{cases}\hskip-14.22636pt\,\,\Rightarrow\,\,\text{iii)}\;\;\frac{\mathrm{I}(-\kappa)^{2}}{\kappa}\sim-\frac{\kappa}{2\ln F(-\kappa)}\,\,.$
Proof.
The proof is analogous to Theorem 3.4, using Lemma 3.1 and 3.3
As a consequence of Theorems 3.4 and 3.5, it holds that if $\ln\bar{F}(\kappa)$ and $\ln F(-\kappa)\in\mathscr{R}_{\theta}$, then
$\frac{\mathrm{I}(\kappa)^{2}}{\kappa}\sim\begin{cases}&\displaystyle-\frac{\kappa}{2\,\ln\bar{F}(\kappa)}\quad\text{as}\;\,\kappa\to+\infty\,\,,\\[14.22636pt] &\displaystyle-\frac{\kappa}{2\,\ln F(\kappa)}\quad\text{as}\;\,\kappa\to-\infty\end{cases}\,\,.$
In many models of practical interest, the asymptotic behaviour of the returns’ distribution is not explicitly known, while the characteristic function of the random variable $\zeta$, defined as
$\phi(\xi):=\mathbb{E}\left[e^{i\xi\zeta}\right]=\int_{-\infty}^{+\infty}e^{i\xi\zeta}\mathrm{d}F(\zeta)\,\,$,
with $\xi$ a complex number with real part $\Re(\xi)$ and imaginary part $\Im(\xi)$, in most cases is explicitly known. We can establish a direct link between the implied volatility formula and the singularity structure of the characteristic function.
Theorem 3.1, p.12 in Lukacs (1972) ensures that if the characteristic function $\phi(\xi)$ “is regular in a neighborhood of the origin then it is also regular in a horizontal strip” bounded by two singularities located on the imaginary axis. We focus on the behaviour of $\phi(\xi)$ near the boundaries of its domain of analyticity. To this end, we introduce the following conditions.
Condition I: $\phi(\xi)$ is analytic in the horizontal strip $\Im(\xi)\in(-\lambda_{+},\lambda_{-})$ with $\lambda_{\pm}>0$ and finite.
Condition II: If Condition I holds and, for some $n\geq 0$,
$(-i)^{n}\phi^{(n)}\left(\pm i\left(\lambda_{\pm}-s\right)\right)\sim s^{-\rho_{\pm}}\,l_{\pm}\left(\frac{1}{s}\right)\quad\text{as $s\to 0^{+}$}\;\;,$
for some $\rho_{\pm}>0$, and $l_{\pm}\in\mathscr{R}_{0}$ and $\phi^{(n)}(u)$ denotes the $n$-th derivative of the function $\phi(\bullet)$, evaluated at $u$.
The latter condition ensures that some derivatives of the characteristic function diverge at $\pm i\lambda_{\pm}$ in a regularly varying manner. As we will see later, the behaviour of the characteristic function in a neighbourhood of these two singularities is crucial, as summarized in Condition II. We state our result in the following Theorem.
Theorem 3.6.
Consider $\lambda_{\pm}>0$ finite. Then, the following implications hold:
$\text{i)}\quad\text{Condition II}\;\Rightarrow\;\text{ii)}\quad\begin{cases}\ln\bar{F}(\kappa)\sim-\lambda_{+}\kappa,\\ \ln F(-\kappa)\sim-\lambda_{-}\kappa\end{cases}\;\;\text{as }\kappa\to+\infty\;\Rightarrow\;\text{iii)}\quad\lim_{\kappa\to\pm\infty}\frac{\mathrm{I}(\kappa)^{2}}{|\kappa|}=\frac{1}{2\lambda_{\pm}}.$
Proof.
We prove that i) $\Rightarrow$ ii). Let us consider the moment generating function of the random variable $\zeta$:
$M(s):=\mathbb{E}\left[e^{s\zeta}\right]=\phi(-is),\quad s\in(-\lambda_{-},\lambda_{+})\,\,.$
Condition I is equivalent to
$M^{(n)}\left(\pm\left(\lambda_{\pm}-s\right)\right)\sim s^{-\rho_{\pm}}\,l_{\pm}\left(\frac{1}{s}\right)\quad\text{as $s\to 0^{+}$}\;\;,$
for some $\rho_{\pm}>0$ and $l_{\pm}\in\mathscr{R}_{0}$, where $M^{(n)}$ denotes the $n$-th derivative of the moment generating function. Indeed
$M^{(n)}\left(\pm\left(\lambda_{\pm}-s\right)\right)=(-i)^{n}\phi^{(n)}\left(\mp i\left(\lambda_{\pm}-s\right)\right)\;\;.$
Thus, we can apply Criterion $1$ from Benaim and Friz (2008), and conclude that
$\begin{cases}\ln\bar{F}(\kappa)\sim-\lambda_{+}\kappa,\\ \ln F(-\kappa)\sim-\lambda_{-}\kappa\end{cases}\quad\text{as}\;\;\kappa\to+\infty\;\;.$
Finally, we prove that ii) $\Rightarrow$ iii). Thanks to ii), $\ln\bar{F}(\kappa)$ and $\ln F(-\kappa)\in\mathscr{R}_{1}$, and we can apply Theorem 3.4 and 3.5
$\lim_{\kappa\to+\infty}\frac{\mathrm{I}(\kappa)^{2}}{\kappa}$ $=-\lim_{\kappa\to+\infty}\frac{\kappa}{2\,\ln\bar{F}(\kappa)}=\frac{1}{2\,\lambda_{+}}\,\,$,
$\lim_{\kappa\to-\infty}\frac{\mathrm{I}(\kappa)^{2}}{\kappa}$ $=-\lim_{\kappa\to-\infty}\frac{\kappa}{2\,\ln{F}(\kappa)}=-\frac{1}{2\,\lambda_{-}}$
Condition II is fairly general and is satisfied by most models commonly used to describe returns in quantitative finance, in which the characteristic function is analytic on the whole horizontal strip $\Im(\xi)\in(-\lambda_{+},\lambda_{-})$. This includes stochastic volatility models (e.g., Heston 1993), Lévy models such as Meixner (see, e.g., Grigelionis, 1999), Tempered Stable, Normal Tempered Stable and Generalized Hyperbolic models (see, e.g., Cont and Tankov, 2003, Ch.4), and additive models (e.g., Carr et al. 2007, Azzone and Baviera 2022). The statemente for Heston model and NTS Lévy models has been proven by Benaim and Friz (2008), while the other cases are discussed in Section 3.1.
For models that generalise the Bachelier framework, [^6] Theorem 3.6 is even more general. Its first implication can be applied to models –commonly used in quantitative finance– defined on log-returns, such as the Heston model or exponential Lévy models (cf., Cont and Tankov 2003, Ch.8). Theorem 3.6 implies that, when at least one derivative of the characteristic function exhibits regularly varying divergence at $\mp i\lambda_{\pm}$, the implied variance grows linearly for large absolute values of moneyness degree $\kappa$, with a slope inversely proportional to the boundaries of the analyticity strip of the characteristic function, as observed in all derivatives markets.
We now prove that the converse also holds. Suppose that the implied variance grows linearly for large absolute values of the moneyness degree $\kappa$. Then the probability distribution has exponentially decaying tails and the characteristic function of returns admits an analyticity strip (Condition I). This is a more general condition than Condition II, and the boundaries of the strip are inversely proportional to the slope of the implied variance.
Theorem 3.7.
Consider $\lambda_{\pm}$ two positive and finite constants. Then, the following implications hold:
$\text{i)}\quad\lim_{\kappa\to\pm\infty}\frac{\mathrm{I}(\kappa)^{2}}{|\kappa|}=\frac{1}{2\lambda_{\pm}}\;\Rightarrow\;\text{ii)}\quad\begin{cases}\ln\bar{F}(\kappa)\sim-\lambda_{+}\kappa,\\ \ln F(-\kappa)\sim-\lambda_{-}\kappa\end{cases}\;\;\text{as }\kappa\to+\infty\;\Rightarrow\;\text{iii)}\;\;\begin{aligned} &\text{Condition I}.\end{aligned}$
Proof.
[…]
$\ln c(\kappa)\sim\frac{\kappa^{2}}{2\,\mathrm{I}(\kappa)^{2}}\sim\lambda_{+}\kappa\quad\text{as }\kappa\to+\infty\,\,$, (17)
thus,
$\ln c(\kappa)\in\mathscr{R}_{1}\,\,.$
Let us observe, from Lemma 3.1, that
$c^{\prime}(\kappa)=-\bar{F}(\kappa)\,\,$,
and that
$(-\ln c(\kappa))^{\prime}=\frac{\bar{F}(k)}{c(k)}\,\,\quad\Rightarrow\quad-\ln(\bar{F}(\kappa))=-\ln\left((-\ln c(\kappa))^{\prime}\right)-\ln(c(\kappa))\,\,.$
Since $\ln c(\kappa)\sim\lambda_{+}\kappa$ as $\kappa\to+\infty$ for (17) and $(-\ln c(\kappa))^{\prime}$ is a positive function of $\kappa$, we can conclude that
$\ln\left((-\ln c(\kappa))^{\prime}\right)=o(\kappa)\quad\text{as }\kappa\to+\infty\quad\Rightarrow\quad-\ln\bar{F}(\kappa)\sim\lambda_{+}\kappa\quad\text{as }\kappa\to+\infty\,\,$,
indeed, applying De l’Hôpital’s rule (whose hypotheses are satisfied thanks to the monotonicity of $\bar{F}(\kappa)$ and $c(\kappa)$,
$\lambda_{+}=\lim_{\kappa\to+\infty}\frac{-\ln c(\kappa)}{\kappa}=\lim_{\kappa\to+\infty}\frac{(-\ln c(\kappa))^{\prime}}{1}=\lim_{\kappa\to+\infty}\frac{\bar{F}(\kappa)}{c(k)}\quad\Rightarrow\quad\lim_{\kappa\to+\infty}\ln(-\ln\bar{F}(\kappa))^{\prime})=\ln(\lambda_{+})\,\,.$
Analogously, we can prove that
$\ln F(-\kappa)\sim\lambda_{-}\kappa\quad\text{as }\kappa\to+\infty\,\,.$
[…]
$\lim_{\kappa\to+\infty}-\frac{\ln\bar{F}(\kappa)}{\kappa}=\lambda_{+}\,\,.$
By definition of the limit, for any arbitrary $a$ satisfying $0<a<\lambda_{+}$, there exists $\kappa_{a}$ such that for all $\kappa>\kappa_{a}$, the inequality $\bar{F}(\kappa)\leq e^{-a\kappa}$ holds. Indeed, choosing $\epsilon>0$ such that $a=\lambda_{+}-\epsilon$, for sufficiently large $\kappa$ we have
$\ln\bar{F}(\kappa)\geq(\lambda_{+}-\epsilon)\,\kappa\,\,$,
which implies
$\bar{F}(\kappa)\leq e^{-(\lambda_{+}-\epsilon)\kappa}=e^{-a\kappa}\,\,.$
Now consider the characteristic function $\phi(\xi)$ for a complex argument $\xi=u+iv$. The existence of the expectation depends on the convergence of the integral involving the real part of the exponent, i.e.,
$\int_{0}^{+\infty}e^{-v\zeta}\mathrm{d}\bar{F}(\zeta)\,\,.$ (18)
Focusing on the positive semi-axis, the integral (18) converges absolutely if $\bar{F}(\kappa)$ decays faster than $e^{v\kappa}$. Using integration by parts and the bound derived above, we obtain:
$\int_{0}^{+\infty}e^{-v\zeta}\mathrm{d}\bar{F}(\zeta)=\left[e^{-v\zeta}\bar{F}(\zeta)\right]_{0}^{+\infty}+v\int_{0}^{+\infty}e^{-v\zeta}\bar{F}(\zeta)\mathrm{d}\zeta<+\infty$
provided that $v>-\lambda_{+}$, since we can always choose $a$ such that $v<a<\lambda_{+}$.
Analogous reasoning applies to the left tail $F(-\kappa)$, ensuring convergence for $v<\lambda_{-}$. Consequently, $\phi(\xi)$ is analytic in the horizontal strip $\Im(\xi)\in(-\lambda_{+},\lambda_{-})$. ∎
The result above is crucial from both a theoretical and a practical perspective. Theorem 3.7 shows that, if the Bachelier implied volatility $\mathrm{I}(\kappa)$ is proportional to the square root of the moneyness degree for large absolute values of $\kappa$, then the model for the underlying returns must belong to the class of models whose characteristic function is analytical in a horizontal strip of the complex plane. This has direct implications for model selection: one may first test whether the implied volatility, which is directly observable in the market, exhibits the behaviour described in point i) of Theorem 3.7, and then select the model accordingly.
In the remaining part of this Section, we show some examples of models for which Condition II holds.
3.1 Examples
In this Section, we show some of the Lévy and additive models already cited above that satisfy Condition II. In particular, we consider: the Meixner process, the Tempered Stable processes and the Generalized Hyperbolic processes.
3.1.1 Meixner
The characteristic function of a Meixner process $\{F_{t}\}_{t\geq 0}$ (with initial condition $F_{0}=0$, see, e.g., Grigelionis 1999, Def.1, p.33) is
$\phi_{F}(\xi):=\left(\frac{\cos\frac{\beta}{2}}{\cosh\frac{\alpha\xi-i\beta}{2}}\right)^{2\delta t}e^{i\xi\mu t}\,\,$,
where $\alpha>0$, $\pi<\beta<\pi$, $\mu\in\mathbb{R}$ and $\delta>0$. The characteristic function of returns $\zeta$ is
$\phi(\xi)=\phi_{F}\left(\frac{\xi}{\sqrt{t}}\right)=\left(\frac{\cos\frac{\beta}{2}}{\cosh\frac{\alpha\xi-i\beta\sqrt{t}}{2\sqrt{t}}}\right)^{2\delta t}e^{i\xi\mu\sqrt{t}}\,\,.$
The characteristic function of $\zeta$ is analytic in the horizontal strip $\Im(\xi)\in(-\lambda_{+},\lambda_{-})$ with
$\lambda_{\pm}=\frac{\pi\mp\beta}{\alpha}\sqrt{t}\,\,.$
With straightforward computations, one can prove that
$(-i)\phi^{\prime}(\mp i(\lambda_{\pm}-s))\sim\sqrt{t}\,\left(\frac{\alpha\,s}{2\sqrt{t}}\right)^{-1-2\delta t}\cos\frac{\beta}{2}\,e^{\pm\lambda_{\pm}\mu\sqrt{t}}\quad\text{as }s\to 0^{+}\,\,$,
thus, Condition II is satisfied with
$\rho_{\pm}=1+2\delta t>0$
$l_{\pm}\left(\frac{1}{s}\right)=\sqrt{t}\,\left(\frac{\alpha}{2\sqrt{t}}\right)^{-1-2\delta t}\cos\frac{\beta}{2}\,e^{\pm\lambda_{\pm}\mu\sqrt{t}}\in\mathscr{R}_{0}\,\,$,
and $n=1$. Furthermore, this result naturally extends to the class of additive processes. Indeed, at any fixed maturity $t$, the time-inhomogeneity of the process does not alter the singularity structure of the characteristic function, ensuring that Condition II remains satisfied.
3.1.2 Tempered Stable
The characteristic function of a Tempered Stable process $\{F_{t}\}_{t\geq 0}$ (with $F_{0}=0$, see, e.g., Cont and Tankov 2003, Prop. 4.2) is
$\phi_{F}(\xi):=\exp\left(t\,C\,\Gamma(-\alpha)\left[(a_{+}-i\xi)^{\alpha}-a_{+}^{\alpha}+(a_{-}+i\xi)^{\alpha}-a_{-}^{\alpha}\right]+i\xi\mu t\right)\,\,$,
where $\alpha\in(0,2)\setminus\{1\}$, $C>0$, and $a_{+},a_{-}>0$. The characteristic function of returns $\zeta$ is
$\phi(\xi)=\phi_{F}\left(\frac{\xi}{\sqrt{t}}\right)=\exp\left(t\,C\,\Gamma(-\alpha)\left[\left(a_{+}-i\frac{\xi}{\sqrt{t}}\right)^{\alpha}-a_{+}^{\alpha}+\left(a_{-}+i\frac{\xi}{\sqrt{t}}\right)^{\alpha}-a_{-}^{\alpha}\right]+i\xi\mu\sqrt{t}\right)\,\,.$
The characteristic function of $\zeta$ is analytic in the horizontal strip $\Im(\xi)\in(-\lambda_{+},\lambda_{-})$ with
$\lambda_{\pm}=a_{\pm}\sqrt{t}\;.$
With straightforward computations, considering the case $0<\alpha<1$, one can prove that
$(-i)\phi^{\prime}(\mp i(\lambda_{\pm}-s))\sim s^{\alpha-1}\sqrt{t}\,C\,\Gamma(1-\alpha)\,t^{-\alpha/2}\phi(\mp i\lambda_{\pm})\quad\text{as }s\to 0^{+}\;,$
thus, Condition II is satisfied with
$\rho_{\pm}=1-\alpha>0$
$l_{\pm}\left(\frac{1}{s}\right)=\displaystyle\sqrt{t}\,C\,\Gamma(1-\alpha)\,t^{-\alpha/2}\,e^{tC\Gamma(-\alpha)[(a_{\mp}+a_{\pm})^{\alpha}-a_{\mp}^{\alpha}-a_{\pm}^{\alpha}]\pm\lambda_{\pm}\mu\sqrt{t}}\in\mathscr{R}_{0}\;,$
and $n=1$. In the case $1<\alpha<2$, the first derivative converges at the boundaries. However, the second derivative diverges. Straightforward computations show that:
$(-i)^{2}\phi^{\prime\prime}(\mp i(\lambda_{\pm}-s))\sim s^{\alpha-2}\sqrt{t}\,C\,\Gamma(2-\alpha)\,t^{-(1+\alpha)/2}\,\phi(\mp i\lambda_{\pm})\quad\text{as }s\to 0^{+}\;,$
thus, Condition II is satisfied with $n=2$ and:
$\rho_{\pm}=2-\alpha>0$
$l_{\pm}\left(\frac{1}{s}\right)=\sqrt{t}\,C\,\Gamma(2-\alpha)\,t^{-(1+\alpha)/2}\,\phi(\mp i\lambda_{\pm})\in\mathscr{R}_{0}\;.$
As observed in the Meixner case, this Condition applies to additive processes as well.
3.1.3 Generalized Hyperbolic models
The characteristic function of a Generalized Hyperbolic process $\{F_{t}\}_{t\geq 0}$ (see, e.g., Cont and Tankov 2003, eq.(4.39), Sec.4.6) is given by
$\phi_{F}(\xi):=e^{i\xi\mu t}\left(\frac{\alpha^{2}-\beta^{2}}{\alpha^{2}-(\beta+i\xi)^{2}}\right)^{\frac{\lambda}{2}}\frac{K_{\lambda}(\delta\sqrt{\alpha^{2}-(\beta+i\xi)^{2}})}{K_{\lambda}(\delta\sqrt{\alpha^{2}-\beta^{2}})}\,\,$,
where $K_{\lambda}(\bullet)$ denotes the modified Bessel function of the second kind (see, e.g., Abramowitz and Stegun 1948, Ch.9, p.374), and the parameters satisfy $\alpha>|\beta|$, $\delta>0$ and $\lambda\in\mathbb{R}$.
The characteristic function of returns $\zeta$ is $\phi(\xi)=\phi_{F}(\xi/\sqrt{t})$ and is analytic in the horizontal strip $\Im(\xi)\in(-\lambda_{+},\lambda_{-})$ with boundaries $\lambda_{\pm}=(\alpha\mp\beta)\sqrt{t}$. The behaviour near the boundary depends on $\lambda$.
The limiting case $\lambda\to 0$, corresponding to the Variance Gamma process, has been discussed in Benaim and Friz (2008).
For $\lambda>0$, the characteristic function diverges at the boundaries; specifically, using the asymptotic relation $K_{\lambda}(z)\sim\frac{1}{2}\Gamma(\lambda)(z/2)^{-\lambda}$ as $z\to 0$ (cf. Abramowitz and Stegun 1948, eq.(9.6.9), Ch.9, p.375) and noting that the argument of the Bessel function scales as $\sqrt{s}$ near the boundary (where $s$ is the distance from the singularity), the function behaves as $\phi(-i(\lambda_{+}-s))\sim s^{-\lambda}$. Thus, Condition II is satisfied with $n=0$ and $\rho_{\pm}=\lambda$.
Conversely, for $\lambda<0$, we define $\nu=-\lambda>0$. The relevant term in the characteristic function takes the form $z^{\nu}K_{\nu}(z)$ with $z\sim\sqrt{s}$. Utilizing the series expansion for the Bessel function of small argument (cf. Abramowitz and Stegun 1948, eq.(9.6.11), Ch.9, p.375), the leading term is constant (implying that the function is bounded), while the dominant singular term scales as $z^{2\nu}\sim s^{\nu}=s^{-\lambda}$. Consequently, the function itself converges, but its derivatives diverge. To satisfy Condition II, one must choose the differentiation order $n$ such that $n>-\lambda$. The $n$-th derivative scales as $s^{-\lambda-n}$, satisfying the condition with a tail decay index $\rho_{\pm}=n+\lambda$. For instance, in the Normal Inverse Gaussian case ($\lambda=-1/2$), we have $n=1$ and $\rho_{\pm}=1/2$.
Similarly, for additive processes of this class, Condition II holds because the time inhomogeneity does not affect the characteristic function’s singularities in the complex plane.
4 Conclusions
In this paper, we have extended the asymptotic analysis of the implied volatility smile from the Black–Scholes setting, as developed by Benaim and Friz (2009), to the Bachelier model, which is particularly relevant in markets where the underlying asset can take negative values.
Using the theory of regular variation, we have derived explicit asymptotic expressions for the Bachelier implied volatility in the wings of the volatility smile, i.e., for large positive or negative moneyness degree. We have proven that there exists a direct link between the tail behaviour of the distribution of the underlying’s returns and the corresponding implied volatility behaviour.
Moreover, in cases where the distribution of returns is not available in closed form but the characteristic function is known, we have shown that our formula can still characterise the asymptotic behaviour of the implied volatility. In particular, we have established that the boundaries of the analyticity strip of the returns’ characteristic function directly determine the behaviour of Bachelier implied volatility in the smile’s wings for most common models in quantitative finance.
Finally, we have shown that if the implied variance grows linearly for large absolute values of the moneyness degree, then the tails of the returns’ probability distribution necessarily exhibit exponential decay and the corresponding characteristic function is analytical in a horizontal strip of the complex plane. Moreover, both the decay rates and the boundaries of the strip are determined by the slope of the implied variance at large moneyness degree.
Acknowledgements
We thank all participants in the Advances in Mathematical Finance Conference in Freiburg for comments and discussions.